OpenStudy (anonymous):

You are given a 2 digit number. 1 times the number you get when you subtract its ones digit from its tens digit is 6; 7 times its tens digit plus 6 times its ones digit is equal to 81. What is the number ?

myininaya (myininaya):

ok let the 10's digit be symbolized by=x let the 1's digit by symbolized by =y x-y=6 7x+6y=81

myininaya (myininaya):

x-y=6 implies x=6+y so 7x+6y=7(6+y)+6y=42+7y+6y=42+13y=81 so 13y=81-42=39 y=39/13=3 x=6+3=9 so the number is 93

OpenStudy (anonymous):

thank you very much